get all points for answering both, get half for answering one or the other

1.PROJECTILE MOTION VELOCITY: A ball is thrown upward from a platform 4.6 m high with a speed of 16 m/s at an angle of 43 o from the horizontal. What is the magnitude of its velocity when it hits the ground?

2. DISPLACEMENT VECTOR: a. A jogger runs directly east for 3 km, then turns and goes northwest for     8 km. He then travels directly south for 2 km. How far is he from the starting point? (km)

       b.In what direction is he from the starting point(measured as an angle counterwise from the              east axis, units are deg)? (Northwest is the direction lying exactly half way between north and          west.)
Answer #1

1)

(v2)^2 = v1^2 - 2g(h)

v2^2 = (16sin(43))^2 - 2*9.8*4.6 = 28.91

v2 = 5.376 m/s


2)

Displacement Vector = 3i - 8cos(45)i + 8sin(45)j - 2j

R = -2.656 i + 3.656 j

How far is he from the starting point? = |R| = sqrt(2.656^2 + 3.656^2) = 4.52 km

b)Direction = arctan(3.656/-2.656) = 180 - 54 = 126 deg counterwise from the east axis